40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Thu Jun 28 23:19:49 2018, johannes, Disha, Notes, Cavity, Gaussian beam mode for the new cavities New_Cavities.pdf
    Reply  Sun Jul 1 23:19:25 2018, johannes, Notes, Cavity, Gaussian beam mode for the new cavities 6x
       Reply  Tue Jul 3 11:58:25 2018, johannes, Notes, Cavity, Gaussian beam mode for the new cavities cavity_support_ring_gen2_15deg.PDFcavity_support_shelf.PDFshelf__contacts.pngcyl__contacts.png
          Reply  Wed Jul 11 22:38:54 2018, johannes, Notes, Cavity, Some assembly required 20180706_212353.jpg20180706_202052.jpg
             Reply  Wed Jul 11 23:27:38 2018, johannes, Notes, Cavity, new breadboard assembly 20180710_143617.jpg20180710_143629.jpg20180710_145112.jpg20180710_163221.jpg20180710_163803.jpg
                Reply  Wed Jul 11 23:51:18 2018, johannes, Notes, Cavity, first light 
                   Reply  Thu Jul 19 23:19:01 2018, johannes, Summary, Cavity, cryostat closed and pumping 20180719_114735.jpg20180717_164839.jpg
                      Reply  Sun Jul 22 00:52:10 2018, johannes, Summary, Cavity, transmission paths set up, mode-matching improved, post-baking ringdowns 6x
Message ID: 2104     Entry time: Thu Jul 19 23:19:01 2018     In reply to: 2103     Reply to this: 2105
Author: johannes 
Type: Summary 
Category: Cavity 
Subject: cryostat closed and pumping 

[Johannes, Disha]

There were more modifications necessary to the breadboard, but today we finally closed the cryostat with the new cavities inside and are in the process of pumping down.


Suspension bracket

The total weight of the breadboard was significantly heavier than before. We removed some spacer mounts but added the lenses + mounts + posts + clamps. As a result it was hanging way too low and would have bottomed on the heat shield. The fix was to alter the mounting brackets, which took some time and required another cycle of cleaning and baking. In the end we turned the brackets upside down cut off the portions that would stick out below the breadboard (Attachment #1). We also noticed that we had to suspend the breadboard even higher than originally planned. The central aperture of the cavity cans inside the cryostat is at 4 inch beam height, same as the viewports, only the breadboard were to rest on the shield bottom. Before, the cavities were a little low in their mounts, which offset that, but the mounts had not taken that into account, so we had to use the lenses to deflect the beams slightly down so we could raise the breadboard and still pass the viewports somewhat centered vertically. This worked out in the end but took extra time. The breadboard was balanced and the cold shield attached to the cold plate, and the cryostat was closed to check the alignment of the beams through the viewports, which checked out.


Mode-matching

The optical table was landed for improving the mode-matching and remains un-floated for the time being. The mode-matching was optimized with the given optics to 87% for the West path and 92% for East. There were some issues with real estate interference on the table for the West, after re-routing it can probably be increased past 90% as well.


Fixing the temperature sensors

When testing the electrical connections the East temperature sensor was acting up, not giving any sensible readings, hinting at a loose connection. This had happened before, and at first I suspected the outside connector of the electrical vacuum feedthrough, because it had been strained during a recent cryostat opening when it was left connected. Opening the connector shell and inspecting the contacts didn't show any broken wires, but upon re-connecting the east sensor seemed fixed? This was only temporary though, and we eventually traced it to a loose connection at the little breakout board of the east sensor directly on the cavity can. The intermittent connection was caused by motion of the wire due to the swinging of the breadboard.


Cryostat closed, new pump line arrangement

We applied a new indium seal and closed the inner volume. The outer can was sealed as well, and the beam alignment quickly checked. We are trying out a new arrangement for the pump line (pics will follow shortly) for which we detached the HiCube's turbo unit from the roughing pump body and seated it on a  breadboard (the old in-vacuum one) that is mounted to the frame that surrounds the optical table. This was done to suppress the vibrations of the vacuum tank from shaking the entire pump line assembly, which was a problem before, particularly when the table was floated. The turbo pump mount is height adjustable, so this arrangement allows for floating the table afterwards. Pumpdown started at ~7pm and by 9 pm the pressure had reached 1 mtorr.

Soon after starting the pumpdown the pump controller displayed the warning code 117, which states that the "pump bottom is getting too hot", referring I believe to the turbo pump (because it was hot), although that was not explicit in the manual. The turbo was struggling to get up to speed, too. I'm not sure why this happened, possibly the pump relied on some heat conduction through the pump body? The turbo fan was working just fine, so I added a fan that blows at the pump from below, and the warning went away after a bit and the turbo eventually sped up to the nominal 90k rpm.

Attachment 1: 20180719_114735.jpg  3.105 MB  Uploaded Fri Jul 20 00:33:24 2018  | Hide | Hide all
20180719_114735.jpg
Attachment 2: 20180717_164839.jpg  3.697 MB  Uploaded Fri Jul 20 00:37:57 2018  | Hide | Hide all
20180717_164839.jpg
ELOG V3.1.3-